# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v1

@@ 4.4 The covariance floor and its strict positivity
$\lim_{B\to\infty}\operatorname{Var}\!\left(\hat{f}_{B}(x)\right)=C_{T}(x),$
so strict positivity of $C_{T}(x)$ establishes an irreducible variance floor at fixed sample size induced by the forest design.
@@
Fix a prediction point $x$. Suppose there exists an index $i^{\star}$ such that $\sigma_{i^{\star}}^{2}>0$ and $\mathbb{P}\!\left(W_{i^{\star}}(x;\theta)>0\right)>0$. Then
$C_{T}(x)>0,\qquad\text{and hence}\qquad\lim_{B\to\infty}\operatorname{Var}\!\left(\hat{f}_{B}(x)\right)>0.$
@@
Theorem 2 shows that repeated reuse of the same training outcomes in local averages at $x$ is sufficient to induce a strictly positive covariance floor. This effect is governed entirely by the resampling design and persists regardless of the number of trees aggregated. The condition $\mathbb{P}\!\bigl(W_{i^{\star}}(x;\theta)>0\bigr)>0$ is satisfied by all standard implementations whenever $x$ lies in the support of the design and at least one observation in its neighborhood has positive conditional variance; moreover, the alignment component provides a second, independent nonnegative contribution to the covariance floor that is not required for strict positivity but reinforces it.
@@ 4.5 Outcome stability of tree-based predictors
Fix a dataset $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a prediction point $x$. Conditional on a realized tree structure, the induced predictor at $x$ can be written in weighted form as
@@ 5.1 The infinite-aggregation target and resolution
Fix a prediction point $x$ and a forest design. The resolution of the infinite-aggregation forest predictor at $x$ is defined by the distribution of the random weight vector $W(x;\theta)=(W_{1}(x;\theta),\ldots,W_{n}(x;\theta))$ induced by the tree-generating mechanism. Equivalently, resolution is characterized by the distribution of the weights appearing in the representation
$f_{\infty}(x)=\sum_{i=1}^{n}\mathbb{E}\!\left[W_{i}(x;\theta)\right]Y_{i}.$
@@
We use the term resolution because it more precisely characterizes the fineness of the forest’s local averaging rule and does not correspond to bias in the conventional parametric sense. Resolution describes the effective granularity of the infinite-aggregation predictor at $x$. Coarser weight distributions average responses over larger sets of observations, yielding increased stability across tree realizations but reduced sensitivity to local variation in the observed outcomes. Finer weight distributions concentrate mass on fewer observations, producing predictors that adapt more sharply to local structure in the realized data but that are more susceptible to dependence across trees through alignment of the induced averaging rules. All forest design parameters affect resolution through their influence on the distribution of the random weight vector $W(x;\theta)$.
@@ 5.2 Training fraction and resolution scaling
$f_{\infty,m}(x)=\mathbb{E}\!\left[T_{\theta}^{(m)}(x)\right].$
@@ 5.3 Structural error pathways
The strict positivity of the covariance floor established in Section 4.4 implies that, even under infinite aggregation, the forest predictor retains design-induced variability determined by the distribution of the weight vector $W(x;\theta)$. While Monte Carlo variability becomes negligible with increasing $B$, the structural component $C_{T}(x)$ reflects how the randomized design allocates weight across training observations and how these allocations co-vary across trees.
Resolution and dependence arise from different aspects of the same design-induced randomness. The infinite-aggregation predictor $f_{\infty}(x)$ is determined by the average weight vector $\mathbb{E}[W(x;\theta)]$, which governs the resolution of the forest at $x$. In contrast, the covariance floor $C_{T}(x)$ depends on the variability and co-movement of the random weights across design realizations. Any design choice that reduces dependence necessarily alters the distribution of these weight vectors, and therefore also affects the resolution of the infinite-aggregation predictor. This relationship is analogous to the classical bias–variance tradeoff. Structural error pathways are thus governed by how the forest design redistributes variability between resolution and dependence, rather than by Monte Carlo variance reduction alone.
@@ 6.3 Aggregation level $B$
For any fixed tree-generating design,
$\mathbb{E}\!\left[\hat{f}_{B}(x)\right]=f_{\infty}(x),\qquad\operatorname{Var}\!\left(\hat{f}_{B}(x)\right)=\frac{1}{B}\,\sigma_{T}^{2}(x)+\frac{B-1}{B}\,C_{T}(x).$
Varying $B$ therefore affects only the Monte Carlo variability arising from finite aggregation. The infinite-aggregation target $f_{\infty}(x)$, the single-tree dispersion $\sigma_{T}^{2}(x)$, and the design-induced covariance $C_{T}(x)$ are properties of the tree-generating mechanism itself and are invariant to $B$.
@@ 6.5 Split-candidate restriction size $q$
where $Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta}]$ denotes the resampling-conditional prediction rule under candidate-set size $q$, and $\theta,\theta^{\prime}$ are independent draws from the tree-generating mechanism. Assume that at each internal node along the path to $x$, the splitting criterion has a unique maximizer over the candidate set almost surely. Then
@@
Resolution at $x$ is governed by the average weight vector $\mathbb{E}[W(x;\theta)]$, which depends on the distribution of the resampling-conditional prediction rule $Z_{\theta}(x)$. Increasing $q$ enlarges the shared candidate information across trees, concentrating the distribution of induced local averaging rules and reducing variability in split selection along the path to $x$. Conversely, decreasing $q$ increases heterogeneity in the induced neighborhoods and broadens the distribution of weight vectors. Thus, larger $q$ produces a more stable and concentrated resolution structure, while smaller $q$ increases diversity of local averaging behavior.
@@ 6.7 Sample size and persistence of structural dependence
$f_{\infty,n}(x)=\mathbb{E}[T_{\theta}^{(n)}(x)],\qquad C_{T,n}(x)=\operatorname{Cov}\!\left(T_{\theta}^{(n)}(x),T_{\theta^{\prime}}^{(n)}(x)\right),$
@@ 7 Discussion
This paper develops a finite-sample, design-based characterization of random forest predictors by treating each tree as a randomized prediction rule acting on fixed data. Under this formulation, all randomness arises from the tree-generating mechanism, and prediction variability is governed by the induced distribution of terminal regions and associated weights. The resulting framework establishes random forests as explicit statistical procedures whose behavior can be analyzed through the same conditioning logic used in classical design-based inference.
@@
A key motivation for this perspective is that random forests differ fundamentally from deterministic prediction procedures. For linear or generalized linear models, the prediction at a fixed point $x$ is an algebraic function of the observed data, and variability arises exclusively from sampling. In random forests, the prediction depends on both the data and the procedural randomization $\theta$, introducing a design-induced variability component that has no analogue in deterministic methods. This additional component decomposes into Monte Carlo noise, which diminishes under sufficient aggregation, and a structural covariance floor $C_{T}(x)$, which does not. The covariance floor is a cost of randomization, paid in exchange for the benefits of ensemble diversity and the ability to trade resolution against stability through hyper-parameters.
@@
The exact variance identity in Theorem 1 makes this separation precise. The Monte Carlo component $\sigma_{T}^{2}(x)/B$ governs seed-to-seed variability and goes to zero as $B$ grows, while $C_{T}(x)$ remains as the irreducible variance floor under infinite aggregation. Decomposing $C_{T}(x)$ via the law of total covariance reveals two structurally distinct mechanisms: observation reuse, whereby shared training outcomes induce correlated predictions, and partition alignment, whereby independently generated trees discover similar local geometry near $x$ and define similar averaging rules even when their training data are disjoint. These mechanisms operate through different features of the tree-generating process and respond to different design parameters – observation reuse to the training fraction $p_{\mathrm{obs}}$, and alignment to the candidate-set size $q$ and the level of split-level randomization.
@@
These results carry direct practical implications. Stabilization of forest predictions across random seeds for tree generation – a standard assessment of sufficient aggregation – reflects only the elimination of the Monte Carlo component and does not indicate the absence or magnitude of $C_{T}(x)$. A forest whose predictions no longer vary across constructions has converged to its infinite-aggregation target $f_{\infty}(x)$, but the precision of that target is governed by the structural covariance between trees. This covariance floor is not detectable through repeated forest construction at fixed $B$; it is a property of the dependence structure among tree-level predictions and can only be assessed by examining their joint behavior within a single forest.
@@
More broadly, existing asymptotic confidence intervals for random forest predictions, including those based on U-statistic and influence-function representations (Mentch and Hooker, 2016; Wager and Athey, 2018), target the sampling variability of the infinite-aggregation functional $f_{\infty}(x)$ under repeated data generation. That is, they quantify how $f_{\infty}(x)$ would vary across new datasets drawn from the same population. Within this sampling-based estimand, these procedures are correctly calibrated. However, they do not incorporate the additional design-induced variability captured by $C_{T}(x)$, which reflects dependence among trees within a fixed dataset and persists under infinite aggregation at fixed sample size. The sampling component decreases at rate $O(n^{-1})$, while $C_{T}(x)$ at fixed design parameters does not necessarily decay at the same rate. When $C_{T}(x)$ is non-negligible relative to the sampling variance, intervals based solely on asymptotic sampling theory quantify sampling uncertainty but may not reflect the total predictive variability of the finite-sample forest procedure.
@@
The present framework complements the existing theoretical literature on random forests. Prior work has established consistency, convergence rates, and asymptotic distributions under increasing sample size, typically by representing forests as weighted nearest-neighbor or U-statistic–type estimators. These results characterize sampling variability but do not isolate the finite-sample dependence induced by the tree-generating mechanism at fixed data. By conditioning on the observed dataset and analyzing the design-induced distribution, the present approach identifies the structural sources of dependence that govern finite-sample behavior and persist under aggregation. Within this formulation, the object of study is the random forest predictor itself, in the same sense that one studies the properties of a fitted generalized linear model independently of how cross-validation is used to assess its performance. Out-of-bag predictions and related resampling-based performance estimates correspond to distinct estimands and should not be conflated with the predictor. Understanding how such estimators inherit and interact with the structural dependence identified here is a natural direction for future work. The covariance floor $C_{T}(x)$ is itself directly estimable from the sample covariance of tree-level predictions within a fitted forest, since these predictions are exchangeable under the design distribution. Decomposing $C_{T}(x)$ into its overlap and alignment components, however, requires the conditional variance quantities $\sigma_{i}^{2}$, which are properties of the data-generating process rather than the randomized design. Developing estimators for these individual components, together with their finite-sample properties, is deferred to subsequent work.
@@ 8 Conclusion
This paper develops a finite-sample, design-based characterization of random forest predictors by viewing each tree as a randomized tree-level prediction rule acting on fixed data. Conditioning on the design yields an exact variance identity that separates aggregation variability from structural dependence induced by the tree-generating mechanism and decomposes this dependence into observation-overlap and partition-alignment components. This framework clarifies the role of randomization in controlling finite-sample prediction variability by weakening structural dependence and reshaping resolution, and provides a principled basis for understanding how forest design parameters govern ensemble behavior beyond aggregation effects alone.
@@ Appendix A Sampling variability of the infinite-forest target
This appendix establishes that the sampling variability of the infinite-aggregation random forest target is generically of order $n^{-1}$ under standard stability conditions.
Let $\mathcal{D}_{n}^{\star}=\{Z_{i}=(X_{i},Y_{i})\}_{i=1}^{n}$ denote an i.i.d. sample from an underlying population, and define the infinite-aggregation forest functional
$f_{\infty}(x;\mathcal{D})=\mathbb{E}_{\Theta}\!\left[\hat{f}_{B}(x;\mathcal{D},\theta)\right].$
Assume that $f_{\infty}(x;\mathcal{D}_{n}^{\star})$ admits an asymptotically linear expansion of the form
@@
so the sampling variability of the infinite-forest target is generically of order $n^{-1}$.

### v2

@@ 4.4 The covariance floor and its strict positivity
$\lim_{B\to\infty}\operatorname{Var}\!\left(\hat{f}_{B}(x)\mid X\right)=C_{T}(x),$
so strict positivity of $C_{T}(x)$ establishes an irreducible variance floor at fixed sample size under the fixed covariate configuration.
@@
Fix a prediction point $x$ and a realized covariate configuration $X$. Suppose there exists an index $i^{\star}$ such that $\sigma_{i^{\star}}^{2}>0$ and
$\mathbb{P}\!\left(W_{i^{\star}}(x;\theta)>0\mid X\right)>0.$
Then
$C_{T}(x)>0,\qquad\text{and hence}\qquad\lim_{B\to\infty}\operatorname{Var}\!\left(\hat{f}_{B}(x)\mid X\right)>0.$
@@
Theorem 2 shows that repeated reuse of the same training outcomes in local averages at $x$ is sufficient to induce a strictly positive covariance floor. This effect is governed entirely by the tree-generating design acting on the fixed covariate configuration and persists regardless of the number of trees aggregated. The condition $\mathbb{P}\!\bigl(W_{i^{\star}}(x;\theta)>0\mid X\bigr)>0$ is satisfied by all standard implementations whenever $x$ lies in the support of the design and at least one observation in its neighborhood has positive conditional variance; moreover, the alignment component provides a second, independent nonnegative contribution to the covariance floor that is not required for strict positivity but reinforces it.
@@ 4.5 Outcome stability of tree-based predictors
Fix a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$ and a prediction point $x$. Conditional on a realized tree structure (that is, conditional on $X$ and a fixed draw of the tree-generating mechanism), the induced predictor at $x$ can be written in weighted form as
@@ 5.1 The infinite-aggregation target and resolution
Fix a prediction point $x$ and a forest design under a realized covariate configuration $X$. The resolution of the infinite-aggregation forest predictor at $x$ is defined by the distribution of the random weight vector $W(x;\theta)=(W_{1}(x;\theta),\ldots,W_{n}(x;\theta))$ induced by the tree-generating mechanism. Equivalently, resolution is characterized by the distribution of the weights appearing in the representation
$f_{\infty}(x)=\sum_{i=1}^{n}\mathbb{E}\!\left[W_{i}(x;\theta)\mid X\right]Y_{i}.$
@@
We use the term resolution because it more precisely characterizes the fineness of the forest’s local averaging rule and does not correspond to bias in the conventional parametric sense. Resolution describes the effective granularity of the infinite-aggregation predictor at $x$ under fixed $X$. Coarser weight distributions average responses over larger sets of observations, yielding increased stability across tree realizations but reduced sensitivity to local variation in the observed outcomes. Finer weight distributions concentrate mass on fewer observations, producing predictors that adapt more sharply to local structure in the realized data but that are more susceptible to dependence across trees through alignment of the induced averaging rules. All forest design parameters affect resolution through their influence on the distribution of the random weight vector $W(x;\theta)$.
@@ 5.2 Training fraction and resolution scaling
$f_{\infty,m}(x)=\mathbb{E}\!\left[T_{\theta}^{(m)}(x)\mid X\right].$
@@ 5.3 Structural error pathways
The strict positivity of the covariance floor established in Theorem 2 implies that, even under infinite aggregation, the forest predictor retains design-induced variability determined by the distribution of the weight vector $W(x;\theta)$. While Monte Carlo variability becomes negligible with increasing $B$, the structural component $C_{T}(x)$ reflects how the randomized design allocates weight across training observations and how these allocations co-vary across trees.
Resolution and dependence arise from different aspects of the same design-induced randomness under fixed $X$. The infinite-aggregation predictor $f_{\infty}(x)=\mathbb{E}[T_{\theta}(x)\mid X]$ is determined by the average weight vector $\mathbb{E}[W(x;\theta)\mid X]$, which governs the resolution of the forest at $x$. In contrast, the covariance floor $C_{T}(x)$ depends on the variability and co-movement of the random weights across design realizations. Any design choice that reduces dependence necessarily alters the distribution of these weight vectors, and therefore also affects the resolution of the infinite-aggregation predictor. This relationship is analogous to the classical bias–variance tradeoff. Structural error pathways are thus governed by how the forest design redistributes variability between resolution and dependence, rather than by Monte Carlo variance reduction alone.
@@ 6.3 Aggregation level $B$
For any fixed tree-generating design and realized covariate configuration $X$,
$\mathbb{E}\!\left[\hat{f}_{B}(x)\mid X\right]=f_{\infty}(x),\qquad\operatorname{Var}\!\left(\hat{f}_{B}(x)\mid X\right)=\frac{1}{B}\,\sigma_{T}^{2}(x)+\frac{B-1}{B}\,C_{T}(x).$
Varying $B$ therefore affects only the Monte Carlo variability arising from finite aggregation. The infinite-aggregation target $f_{\infty}(x)$, the single-tree dispersion $\sigma_{T}^{2}(x)$, and the design-induced covariance $C_{T}(x)$ are properties of the tree-generating mechanism itself under fixed $X$ and are invariant to $B$.
@@ 6.5 Split-candidate restriction size $q$
where $Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta},X]$ denotes the resampling-conditional prediction rule under candidate-set size $q$, and $\theta,\theta^{\prime}$ are independent draws from the tree-generating mechanism. Assume that at each internal node along the path to $x$, the splitting criterion has a unique maximizer over the candidate set almost surely. Then
@@
Resolution at $x$ is governed by the average weight vector $\mathbb{E}[W(x;\theta)\mid X]$, which depends on the distribution of the resampling-conditional prediction rule $Z_{\theta}(x)$. Increasing $q$ enlarges the shared candidate information across trees, concentrating the distribution of induced local averaging rules and reducing variability in split selection along the path to $x$. Conversely, decreasing $q$ increases heterogeneity in the induced neighborhoods and broadens the distribution of weight vectors. Thus, larger $q$ produces a more stable and concentrated resolution structure, while smaller $q$ increases diversity of local averaging behavior.
@@ 6.7 Sample size and persistence of structural dependence
$f_{\infty,n}(x)=\mathbb{E}[T_{\theta}^{(n)}(x)\mid X],\qquad C_{T,n}(x)=\operatorname{Cov}\!\left(T_{\theta}^{(n)}(x),T_{\theta^{\prime}}^{(n)}(x)\mid X\right),$
@@ 7 Discussion
This paper develops a finite-sample, design-based characterization of random forest predictors by treating each tree as a randomized prediction rule acting on a fixed covariate configuration. Conditional on $X$, randomness arises jointly from the outcome realization $(Y\mid X)$ and the tree-generating mechanism, and prediction variability is governed by the induced distribution of terminal regions and associated weights. The resulting framework establishes random forests as explicit statistical procedures whose behavior can be analyzed through the same conditioning logic used in classical fixed-design inference.
@@
A key motivation for this perspective is that random forests differ fundamentally from deterministic prediction procedures. For linear or generalized linear models, the prediction at a fixed point $x$ is an algebraic function of the observed data, and variability arises exclusively from sampling. In random forests, the prediction depends on both the data and the procedural randomization $\theta$, introducing a design-induced variability component that has no analogue in deterministic methods. Conditional on $X$, this additional component decomposes into Monte Carlo noise, which diminishes under sufficient aggregation, and a structural covariance floor $C_{T}(x)$, which does not. The covariance floor is a cost of randomization, paid in exchange for the benefits of ensemble diversity and the ability to trade resolution against stability through hyper-parameters.
@@
The exact variance identity in Theorem 1 makes this separation precise. The Monte Carlo component $\sigma_{T}^{2}(x)/B$ governs seed-to-seed variability and goes to zero as $B$ grows, while $C_{T}(x)$ remains as the irreducible variance floor under infinite aggregation conditional on $X$. Decomposing $C_{T}(x)$ via the law of total covariance reveals two structurally distinct mechanisms: observation reuse, whereby shared training outcomes induce correlated predictions, and partition alignment, whereby independently generated trees discover similar local geometry near $x$ and define similar averaging rules even when their training data are disjoint. These mechanisms operate through different features of the tree-generating process and respond to different design parameters – observation reuse to the training fraction $p_{\mathrm{obs}}$, and alignment to the candidate-set size $q$ and the level of split-level randomization.
@@
These results carry direct practical implications. Stabilization of forest predictions across random seeds for tree generation – a standard assessment of sufficient aggregation – reflects only the elimination of the Monte Carlo component and does not indicate the absence or magnitude of $C_{T}(x)$. A forest whose predictions no longer vary across constructions has converged to its infinite-aggregation target $f_{\infty}(x)=\mathbb{E}[T_{\theta}(x)\mid X]$, but the precision of that target is governed by the structural covariance between trees. This covariance floor is not detectable through repeated forest construction at fixed $B$; it is a property of the dependence structure among tree-level predictions and can only be assessed by examining their joint behavior within a single forest.
@@
More broadly, existing asymptotic confidence intervals for random forest predictions, including those based on U-statistic and influence-function representations (Mentch and Hooker, 2016; Wager and Athey, 2018), target the sampling variability of the infinite-aggregation functional $f_{\infty}(x)$ under repeated data generation. That is, they quantify how $f_{\infty}(x)$ would vary across new datasets drawn from the same population. Within this sampling-based estimand, these procedures are correctly calibrated. However, they do not incorporate the additional design-induced variability captured by $C_{T}(x)$, which reflects dependence among trees within a fixed dataset and persists under infinite aggregation at fixed sample size. The sampling component decreases at rate $O(n^{-1})$ under standard stability conditions, while $C_{T}(x)$ at fixed design parameters need not decay at the same rate. When $C_{T}(x)$ is non-negligible relative to the sampling variance, intervals based solely on asymptotic sampling theory quantify sampling uncertainty but may not reflect the total predictive variability of the finite-sample forest procedure.
@@
The present framework complements the existing theoretical literature on random forests. Prior work has established consistency, convergence rates, and asymptotic distributions under increasing sample size, typically by representing forests as weighted nearest-neighbor or U-statistic–type estimators. These results characterize sampling variability but do not isolate the finite-sample dependence induced by the tree-generating mechanism at fixed data. By conditioning on the observed covariate configuration and analyzing the design-induced distribution, the present approach identifies the structural sources of dependence that govern finite-sample behavior and persist under aggregation. Within this formulation, the object of study is the random forest predictor itself, in the same sense that one studies the properties of a fitted generalized linear model independently of how cross-validation is used to assess its performance. Out-of-bag predictions and related resampling-based performance estimates correspond to distinct estimands and should not be conflated with the predictor. Understanding how such estimators inherit and interact with the structural dependence identified here is a natural direction for future work. The covariance floor $C_{T}(x)$ is itself directly estimable from the sample covariance of tree-level predictions within a fitted forest, since these predictions are exchangeable under the design distribution conditional on $X$. Decomposing $C_{T}(x)$ into its overlap and alignment components, however, requires the conditional variance quantities $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$, which are properties of the outcome model rather than the randomized design. Developing estimators for these individual components, together with their finite-sample properties, is deferred to subsequent work.
@@ 8 Conclusion
This paper develops a finite-sample, design-based characterization of random forest predictors by viewing each tree as a randomized tree-level prediction rule acting on a fixed covariate configuration. Conditioning on $X$ yields an exact variance identity that separates aggregation variability from structural dependence induced by the tree-generating mechanism and decomposes this dependence into observation-overlap and partition-alignment components. This framework clarifies the role of randomization in controlling finite-sample prediction variability by weakening structural dependence and reshaping resolution, and provides a principled basis for understanding how forest design parameters govern ensemble behavior beyond aggregation effects alone.
@@ Appendix A Sampling variability of the infinite-forest target
This appendix studies the second term in the unconditional variance decomposition presented in Section 2.1,
$\operatorname{Var}\!\left(\mathbb{E}\!\bigl[\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid X\bigr]\right),$
which captures how the forest predictor’s conditional mean varies across repeated draws of the covariate configuration from the population. For large $B$, $\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)$ approximates its infinite-aggregation limit $f_{\infty}(x;\mathcal{D}_{n}^{\star})=\mathbb{E}_{\theta}[T_{\theta}(x;\mathcal{D}_{n}^{\star})]$, and the conditional expectation $\mathbb{E}[\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid X]$ converges to $\mathbb{E}[f_{\infty}(x;\mathcal{D}_{n}^{\star})\mid X]$. In the infinite-aggregation limit, the second term therefore reduces to
$\operatorname{Var}\!\left(\mathbb{E}\!\bigl[f_{\infty}(x;\mathcal{D}_{n}^{\star})\mid X\bigr]\right),$
which measures how the design-based target of the forest changes across repeated draws of $X$ from the population. Because $f_{\infty}(x;\mathcal{D}_{n}^{\star})$ depends on both $X$ and $Y$, the inner expectation averages over $Y\mid X$ and the remaining variability is driven by sampling of the covariate configuration itself.
To bound this term, we establish a rate for the stronger quantity $\operatorname{Var}(f_{\infty}(x;\mathcal{D}_{n}^{\star}))$, which provides an upper bound since $\operatorname{Var}(\mathbb{E}[f_{\infty}\mid X])\leq\operatorname{Var}(f_{\infty})$ by the law of total variance.
Let $\mathcal{D}_{n}^{\star}=\{Z_{i}=(X_{i},Y_{i})\}_{i=1}^{n}$ denote an i.i.d. sample from an underlying population. Assume that $f_{\infty}(x;\mathcal{D}_{n}^{\star})$ admits an asymptotically linear expansion of the form
@@
and therefore
$\operatorname{Var}\!\left(\mathbb{E}\!\bigl[f_{\infty}(x;\mathcal{D}_{n}^{\star})\mid X\bigr]\right)\leq\frac{\operatorname{Var}(\psi_{x}(Z_{1}))}{n}+o(n^{-1}).$
The sampling variability of the infinite-forest target is thus generically of order $O(n^{-1})$, justifying the focus in the main text on the design-induced variability captured by the first term $\mathbb{E}[\operatorname{Var}(\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid X)]$, which contains the structural covariance floor $C_{T}(x)$ that persists under infinite aggregation at fixed sample size.